### v2

# Shaken or Stirred? An Analysis of MetaFormer’s Token Mixing for Medical Imaging

Abstract
The generalization of the Transformer architecture via MetaFormer has reshaped our understanding of its success in computer vision. By replacing self-attention with simpler token mixers, MetaFormer provides strong baselines for vision tasks. However, while extensively studied on natural image datasets, its use in medical imaging remains scarce, and existing works rarely compare different token mixers, potentially overlooking more suitable designs choices. In this work, we present the first comprehensive study of token mixers for medical imaging. We systematically analyze pooling-, convolution-, and attention-based token mixers within the MetaFormer architecture on image classification (global prediction task) and semantic segmentation (dense prediction task). Our evaluation spans eight datasets covering diverse modalities and common challenges in the medical domain. Given the prevalence of pretraining from natural images to mitigate medical data scarcity, we also examine transferring pretrained weights to new token mixers. Our results show that, for classification, low-complexity token mixers (e.g. grouped convolution or pooling) are sufficient, aligning with findings on natural images. Pretrained weights remain useful despite the domain gap introduced by the new token mixer. For segmentation, we find that the local inductive bias of convolutional token mixers is essential. Grouped convolutions emerge as the preferred choice, as they reduce runtime and parameter count compared to standard convolutions, while the MetaFormer’s channel-MLPs already provide the necessary cross-channel interactions. Our code is available at https://github.com/multimodallearning/MetaFormerMedImaging/tree/clean_code.
keywords:
MetaFormer, Local Self-Attention, Medical Imaging, Fine-Tuning, Large Kernels, Classification, Segmentation
doi:
10.59275/j.melba.2024-AAAA
volume: 3
CNN
Convolutional Neural Network
NLP
Natural Language Processing
ViT
Vision Transformer
MLP
Multi-Layer Perceptron
SAR
Synthetic Aperture Radar
CT
Computer Tomography
MRI
Magnetic Resonance Imaging
FLOPs
Floating Point Operations
AUC
Area under the Receiver Operating Characteristic curve
DSC
Dice Similarity Score
DETR
Detection Transformer
GAP
Global Average Pooling
1 Introduction
[…]
Over the last ten years, Deep Learning and especially Convolutional Neural Networks have set the state-of-the-art in various tasks of medical image processing like classification (Anwar et al., 2018), segmentation (Isensee et al., 2021) and detection (Fernando et al., 2021). Following its success in Natural Language Processing (NLP(Devlin et al., 2019; Radford et al., 2019), self-attention (Vaswani et al., 2017) has soon emerged as an alternative to convolution. The Transformer architecture has been successfully adapted to image classification (Dosovitskiy et al., 2021), segmentation (Xie et al., 2021) and detection (Carion et al., 2020). Further adaptation of the Transformer architecture targeted specific requirements for vision tasks. This includes for example a hierarchical pyramid-like architecture like it can be found in CNNs (Wang et al., 2021; Liu et al., 2021), the limitation of its receptive field to the direct neighbourhood for detection (Zhu et al., 2020), or tackling characteristics of self-attention that hinder the usage of Vision Transformers as the backbone for further vision tasks (Darcet et al., 2023).
The property of self-attention having a global receptive field is a doubled-edged sword in image processing where large input sizes are common. On the one hand, the number of trainable parameters of the self-attention operator is independent of the input size, and it enables capturing of large context. On the other hand, since every pixel attends to each other pixel in an input, its computational complexity grows quadratically with the resolution of 2D images, making it inefficient for high resolutions. In contrast, convolution limits its receptive field to the direct neighbourhood of the pixel and thus provides a strong local bias and translation-equivariance, two properties fitting the domain of images. Combining both paradigms, many works have also limited the receptive field of self-attention, introducing local self-attention. Here, the implementations can be divided into two groups, splitting the input into patches (Vaswani et al., 2021) or limiting the influence of pixels outside the direct neighbourhood by applying a (soft) weighting during the self-attention operation (d’Ascoli et al., 2021). For both local and global attention, implementations were introduced (Pan et al., 2023; Dao et al., 2022). Further, recent work investigates into the principle of filter separation, bringing state space models to computer vision within the MAMBA architecture (Zhu et al., 2024). Lately, frameworks like PyTorch’s flexattention enables the easy implementation of various attention patterns.
There also exist efforts (e.g. classification (Liu et al., 2022) and segmentation (Lee et al., 2022; Roy et al., 2023)) to enable large receptive fields for CNNs without equipping the model with too many learnable parameters.
While Tolstikhin et al. (2021) demonstrate that replacing the self-attention within the ViT with Multi-Layer Perceptron (MLP) yields comparable performance (Dosovitskiy et al., 2021; Vaswani et al., 2021), the MetaFormer framework (Yu et al., 2022) takes this a step further by abstracting the general transformer architecture into two parts: The token mixer (e.g. self-attention) learns to aggregate spatial information and the channel-MLP extracts new features with a point-wise operation. Within this MetaFormer perspective, using grouped convolution as the token mixer naturally draws a parallel to ConvNeXt (Liu et al., 2022), where depthwise convolution fulfils an analogous role in aggregating spatial information. However, using pooling as a simple, non-parameterised token mixer, the MetaFormer outperforms ViT on ImageNet. This implies that the channel-MLPs provide the main contribution to the transformer’s capabilities in computer vision, rather than the token mixer. In their following work, Yu et al. (2023) provide strong results (ImageNet top-1 accuracy of over $80\,\%$) using no token mixing at all. Still, equipping the token mixers with some learnable parameters was found beneficial (see Yu et al. (2022), Tab. 5).
The general formulation of the MetaFormer easily allows for the deployment of hybrid architectures, combining different token mixers (e.g. using pooling as an efficient token mixer in the first stages with high-resolution images and global self-attention in latter ones). Since then, MetaFormer has been implemented in various applications and extended with specialised token mixers (see Sec. 2).
1.1 Contribution
Most existing applications of the MetaFormer lie outside the medical domain, and those within typically employ it without comparing different token mixers – potentially overlooking more suitable choices. This work presents the first comprehensive study of token mixers for medical imaging, analysing their strengths, weaknesses, and complexity in terms of trainable parameters and runtime. We systematically compare six token mixers, varying the kernel size for kernel-based mixers such as convolution and pooling. Our evaluation covers eight datasets, including seven from the medical domain and one natural image benchmark. We investigate two different tasks: classification as a global prediction task (five datasets) and semantic segmentation as a dense prediction task (three datasets, covering small, medium, and large resolutions). The datasets represent diverse modalities, including histology, microscopy, and radiography. With this, we chose to expose the different token mixers to key medical imaging challenges such as large homogeneous regions in abdominal CT and long-range context requirements in pathology. We summarise our key findings as follows.
  • Datasets containing a single dominant object in the image favour more locality and thus smaller kernel sizes. Conversely, datasets with a larger region of interest tend to profit from incorporating a broader spatial context through larger kernels.
  • For classification, the MetaFormer trained from scratch with no token mixing already yields very competitive performance. The results are best when using low-complexity token mixers.
  • When pretrained weights are available, exchanging token mixers introduces a domain gap. But pretrained weights of global attention can still be successfully utilized in local attention. This provides a more efficient alternative to global attention when dealing with large input resolutions.
  • For segmentation, we observe that a strong local inductive bias is crucial, making grouped convolutional token mixers the most effective choice.
2 Related Works
With the right choice of token mixer, the MetaFormer can be a light-weighted model, enabling its use for real-time applications like ship signature classification in Synthetic Aperture Radar (SAR) (Zhu et al., 2023) or in autonomous driving (Shun et al., 2024), which deploys additional token pruning for reduced runtime complexity. For super resolution of meteorological satellite cloud images, Wang et al. (2023) introduce a new token mixer combining gated convolution and channel-space attention. With the extension of temporal pooling, the MetaFormer has been applied on time series, capturing the structural and temporal relationship in human motion prediction for human-robot interaction (Xu et al., 2023). A novel idea of token mixing moves this operation to the Fourier spectrum, enabling efficient global token mixing without limited receptive field, but adding the cost of the Fourier transformation and its inverse. With balancing the low- and high-pass characteristics via spectral pooling aggregation (Yun et al., 2023), the performance on ImageNet is improved by $3\,\%$ accuracy compared to the PoolFormer (MetaFormer with average pooling as token mixer) (Yu et al., 2022). In Tatsunami and Taki (2024), this approach is extended, so coefficients of a dynamic filter bank are estimated by an MLP based on the input. While the original work (Yu et al., 2022) employs the MetaFormer itself as a backbone for semantic segmentation, MetaSeg (Kang et al., 2024) extends its principles to a decoder and introduces channel reduction attention for gathering global context within the decoder.
In the medical domain, MetaFormer mostly gets adapted for semantic segmentation. For polyp segmentation, MetaFormer has been paired with a parallel path comprising convolutional units with residual and dilated paths to compensate for MetaFormer’s weak low-level features (Lee and Yoo, 2024). Another work on polyp segmentation (Trinh et al., 2023) utilizes the MetaFormer as the encoder in a U-Net-like architecture and extend the decoder’s dilated convolutions with learnable multiscale upsampling. In volumetric brain tumour segmentation, MetaFormer has been used as an auxiliary decoder, combining local–global pooling with depthwise convolution as the token mixer. To improve efficiency, it has been adapted with involution and a squeeze-and-excitation module in place of channel-MLP (Dwika Hefni Al-Fahsi et al., 2024). Similarly, MAMBA’s structured visual space model has been integrated into MetaFormer for image restoration of brain Magnetic Resonance Imaging (MRI) and abdominal Computer Tomography (CT) (Chen et al., 2024). For pancreas tissue detection, deformable convolutions have been employed as token mixers, dynamically adapting the receptive field (Kurami et al., 2024). To the best of our knowledge, this is the first comprehensive study of different standard token mixers in the MetaFormer on medical image analysis.
3 Methods
Figure 1: Schema of the MetaFormer block illustrating different token mixer options employed across stages of the MetaFormer architecture (cf. Fig. 2(b)).
((a)) FLOPs of the token mixers across the four MetaFormer stages (stage index: $C\times N$) for an input of size $3\times 768^{2}$.
((b)) Extension of MetaFormer for classification by collapsing the final feature map to a vector and segmentation employing the SegFormer decoder Xie et al. (2021) with number of classes $|\Omega|$ and latent dimension $C=256$ for segmentation.
Figure 2: Visualization of FLOPs for different token mixers (2(a)) and schema of MetaFormer’s architecture (2(b)).
In a typical medical image processing task, we are given an input image $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$ and aim to learn a function $f(\mathbf{I})$ that projects the image into a predefined solution space. In image classification, a vector in this solution space is described as a global vector $\mathbf{s}\in\mathbb{R}^{|\Omega|}$ holding the likelihood $p(\mathbf{I}|\omega_{n})$ for each class $\omega\in\Omega$. For semantic segmentation, the classification vector is defined for every pixel $\mathbf{p}\in\mathbf{I}$. Thus, we obtain a segmentation mask $\mathbf{S}\in\mathbb{R}^{|\Omega|\times H\times W}$ holding the likelihoods for every pixel $p(\mathbf{p}|\omega)$ to belong to class $\omega$. When approximating the optimal $f$ in a given task with deep learning, we used to implement $f=f_{0}\circ f_{1}\circ\dots\circ f_{L}$ as a nested chain of functions $f_{l}$, where each $\{f_{l}\}_{l=0}^{L}$ represent a layer in a neural network. In the past ten years, the favoured operator in image processing for $f_{l}$ has been convolution (Krizhevsky et al., 2012; He et al., 2016; Liu et al., 2022). After their initial success in natural language processing tasks, Transformers with their self-attention operator have also emerged as an alternative to convolutional networks (Dosovitskiy et al., 2021; Xie et al., 2021; Carion et al., 2020). However, while ViTs achieve state-of-the-art performance in many image tasks, it remains unclear, which architectural components drive their success. To address this, MetaFormer (Yu et al., 2022, 2023) generalizes the Transformer architecture by replacing self-attention with operations better suited to vision.
3.1 MetaFormer Architecture
The MetaFormer (Yu et al., 2022) block represents a generalized version of the Transformer in that it defines token mixing as a variable operation, of which self-attention is only one of many possible operations. The other structures of the Transformers (e.g. channel MLP or residual connections) have been proven to be also crucial for computer vision (Yu et al., 2022) and are therefore left untouched. Fig. 1 shows the structure of the MetaFormer’s block. Within its multi-stage architecture, each employing multiple MetaFormer blocks, the MetaFormer adopts the hierarchical approach of CNNs (He et al., 2016; Krizhevsky et al., 2012) of halving the spatial resolution and doubling the feature channels at the beginning of each stage using a convolution layer (see Fig. 2(b)). The first stage is an exception with a downsampling factor of four. The number of blocks per stage is a hyperparameter to modify model capacity, similar to the ResNet (He et al., 2016). With this, the MetaFormer enables feature extraction on different resolution scales, which stands in contrast to the original ViT (Dosovitskiy et al., 2021), where the image is patchified and then fed as a sequence of patches through the network without deviating from the original resolution.
3.2 Different Token Mixers
Following MetaFormer (Yu et al., 2022, 2023), we investigate commonly used operations as token mixers denoted by $t(x)$. We compare identity, average pooling, convolution, grouped convolution, self-attention, and local self-attention, all defined in the following paragraphs. The choice of token mixer introduces fundamental differences in properties such as the number of trainable parameters and runtime complexity. Moreover, each token mixer carries distinct inductive biases regarding locality, making them more appropriate for particular imaging domains (e.g., those requiring strong local interactions, long-range context, or texture-based representations).
The computational complexity and trainable parameter count of the investigated token mixers including the channel-MLP of the MetaFormer block are summarized in Tab. 1. We state the complexity in terms of the input size: number of channels $C$, height $H$, and width $W$ (for shorter notation we define $N=HW$). For local operators, we define their kernel or pooling size $K$, where we only consider kernels of size $(K\times K)$ containing $K^{2}$ elements, since it is the most common choice. Note that a MetaFormer block does not change the spatial resolution or number of channels and, therefore, we use the same variables for in- and output sizes.
To define the following operations, we introduce the indices $c,h,w\in\mathbb{N}$ iterating over the $C$ channels, the height $H$ and width $W$. We define the pixel value at the index $(c,h,w)$ as $\mathbf{I}(c,h,w)=p\in\mathbb{R}$.
3.2.1 Identity
Although strictly speaking, the identity provides no token mixing, we include it as a lower baseline to investigate which benefits token mixing provides on its own.
$t_{\texttt{id}}\left(\mathbf{I},c,h,w\right)=\mathbf{I}(c,h,w)$ (1)
The identity operation introduces neither additional computational cost nor additional parameters to the MetaFormer architecture.
3.2.2 Average Pooling
As a next step, we introduce spatial average pooling. This operation performs basic token mixing and adds a local inductive bias beneficial for image processing, while requiring no additional parameters. The average is computed within a pool of size $K^{2}$:
$t_{\texttt{avg}}\left(\mathbf{I},c,h,w\right)=\frac{1}{K^{2}}\sum_{i=-K}^{K}\sum_{j=-K}^{K}\mathbf{I}(c,h+i,w+j)$ (2)
With this, its computational complexity is $\mathcal{O}(NK^{2}C)$.
3.2.3 Convolution
Convolution provides with its locality a favourable inductive bias for data on regular grids like images with highly correlated features in neighbourhoods. By optimizing its kernel parameters $\mathbf{K}\in\mathbb{R}^{C\times C\times K\times K}$ by gradient descent (LeCun et al., 1989), it has been responsible for the breakthrough of Deep Learning in computer vision, providing a higher efficiency through its sparsity and reuse of parameter when compared to previously standard fully connected models. All this makes convolution a natural choice as a token mixer in image-based tasks.
$t_{\texttt{conv}}$ $\left(\mathbf{I},\mathbf{K},c_{\texttt{out}},h,w\right)$ (3)
$=\sum_{c=1}^{C}\sum_{i=-K}^{K}\sum_{j=-K}^{K}\mathbf{I}(c,i,j)\mathbf{K}(c_{\texttt{out}},c,h+i,w+j)$
Due to its connectivity over the channels, convolution has a high computational complexity of $\mathcal{O}(NK^{2}C^{2})$ and uses $K^{2}C^{2}$ parameters, mostly limiting the choice of kernel size to $K=3$ for all but the first layers (Krizhevsky et al., 2012; He et al., 2016).
3.2.4 Grouped Convolution
To overcome the high complexity of convolution, grouped convolution can be used (Krizhevsky et al., 2012). Using $G$ filter groups, an output channel no longer depends on all input channels but on a subset of $\frac{C}{G}$. Thus, the depth connectivity can be reduced ($G<C$) or omitted ($G=C$). Note that $G$ has to be chosen so that $C$ is an integer multiple of $G$. To show the extrema, we chose $G=C$, shrinking the kernel weight matrix to $\mathbf{K}\in\mathbb{R}^{C\times K\times K}$ and thus omit the sum over the channel (compare Eq. 3)
$t_{\texttt{group\_conv}}$ $\left(\mathbf{I},\mathbf{K},c,h,w\right)$ (4)
$=\sum_{i=-K}^{K}\sum_{j=-K}^{K}\mathbf{I}(c,i,j)\mathbf{K}(c,h+i,w+j)$
This reduces complexity to $\mathcal{O}(NK^{2}C)$ and $K^{2}C$ parameters. Since channels are processed independently, grouped convolution with $G=C$ lacks channel cross-talk. However, as each MetaFormer block includes a channel-MLP (Fig. 1) introducing cross-talk, we argue this limitation is negligible. This design parallels the depthwise separable convolution of MobileNet (Howard et al., 2017).
3.2.5 Self-Attention
With Self-Attention, Transformers have set a new state-of-the-art in natural language processing (Vaswani et al., 2017) and have also been well adapted to image processing tasks (Dosovitskiy et al., 2021; Xie et al., 2021; Carion et al., 2020). However, Yu et al. (2022) provide evidence that the self-attention mechanism may not be the most important part of a successful model architecture for computer vision. To test this observation in the medical imaging domain, we include self-attention as a baseline for token mixing.
Let $\mathbf{p}_{r}=\mathbf{I}(h,w)=\mathbf{I}(\lfloor\tfrac{r}{W}\rfloor,r\mod W)\in\mathbb{R}^{C}$ define the current pixel’s feature vector with $r=h\cdot W+w\leq N$ as the ravelled, flat index of $(h,w)$. For self-attention as token mixer $t_{\texttt{global\_attn}}$, each pixel $\mathbf{p}_{r}\in\mathbb{R}^{C}$ is transformed into the three roles: key $\mathbf{k}_{r}=\mathbf{W}_{k}\mathbf{p}_{r}$, value $\mathbf{v}_{r}=\mathbf{W}_{v}\mathbf{p}_{r}$, and query $\mathbf{q}_{r}=\mathbf{W}_{q}\mathbf{p}_{r}$, where each $\mathbf{W}\in\mathbb{R}^{C\times C}$ describes a learnable weight matrix. The transformed value of $\mathbf{p}_{r}$ is obtained by
$\mathbf{z}_{m,r}$ $=\sum_{s=1}^{N}\underset{s=1,\dots,N}{\texttt{softmax}}\left(\frac{\mathbf{q}_{m,r}^{\intercal}\mathbf{k}_{m,s}}{\sqrt{D}}\right)\cdot\mathbf{v}_{m,s}$ (5)
$\mathbf{z}_{r}$ $=\mathbf{W}_{u}\texttt{concat}(\{z_{m,r}\}_{m=1}^{M}),$ (6)
where $\mathbf{k},\mathbf{v}$, and $\mathbf{q}$ is split up into $M$ vectors of size $D$, forming attention heads $m=1,\dots,M$, allowing to learn different attention patterns. Their results are finally unified to $\mathbf{z}_{r}$ by $\mathbf{W}_{u}\in\mathbb{R}^{C\times C}$. In our experiments, we adapt the number of heads $M=\frac{C}{16}$ dynamically for each stage, dividing the number of channels $C$ by 16. Since self-attention is permutation-invariant, but images have an informative locality structure, we add positional embedding following Yu et al. (2022). While self-attention’s parameter count $4C^{2}$ is independent of the amount of processed pixels, its computational complexity quadratically grows with their amount $\mathcal{O}(N^{2}C)$. This significantly limits the applicability of self-attention for large input resolutions.
3.2.6 Local Self-Attention
To break up the quadratic runtime complexity and to infuse self-attention with the same advantages of convolution’s locality (e.g. inductive bias and translation equivariance), we limit the number of keys a query can attend to its direct neighbourhood. Let $\mathcal{N}$ define a function that returns the neighbouring pixel coordinates of a given location:
$\mathcal{N}_{K}(h,w)=\left\{\begin{aligned} i,j\in\mathbb{N}_{0}:\;&(|i-h|<\tfrac{K}{2})\\ &\land(|j-w|<\tfrac{K}{2})\end{aligned}\right\},$ (7)
where $K$ describes the kernel size in height and width, since we limit ourselves to quadratic kernels. We equivalently denote the neighbourhood function as $\mathcal{N}_{K}(r)$ with the flat index $r$. With it, we extend Eq. (6) to construct $t_{\texttt{local\_attn}}$ as follows:
$\mathbf{z}_{m,r}=\sum_{s\in\mathcal{N}_{K}(r)}\underset{s\in\mathcal{N}_{K}(r)}{\texttt{softmax}}\left(\frac{\mathbf{q}_{m,r}^{\intercal}\mathbf{k}_{m,s}}{\sqrt{D}}\right)\cdot\mathbf{v}_{m,s}.$ (8)
We implement $\mathcal{N}_{K}(r)$’s attention pattern using PyTorch’s flexattention framework111pytorch.org/blog/flexattention. By restricting the receptive field of $\mathbf{q}$, local self-attention reduces complexity to $\mathcal{O}(NK^{2}C)$ for typical resolutions ($K^{2}\ll N=HW$) compared to global self-attention (cf. Fig. 2(a)), while keeping the parameter count unchanged at $4C^{2}$.
Table 1: Runtime complexity in $\mathcal{O}$ notation, Floating Point Operations (FLOPs), and trainable parameter count of different token mixers in the MetaFormer block (plus channel-MLP) regarding the kernel size $K$, number of in- and output channels $C$, and spatial input size $N=H\cdot W$. Note that in MetaFormer, no token mixer changes the spatial resolution or number of feature channels. Similarly, we only consider kernels of shape $(K\times K)$ comprising $K^{2}$ elements. FLOPs formulas are constructed ignoring technical optimisations in hard- and software.
Token Mixer Runtime Complexity FLOPs formula # Parameters
Identity $\mathcal{O}(NC^{2})$ $NC^{2}$ $C^{2}$
Average Pooling $\mathcal{O}(NK^{2}C+NC^{2})$ $NK^{2}C+NC^{2}$ $C^{2}$
Grouped Convolution $\mathcal{O}(NK^{2}C+NC^{2})$ $N2K^{2}C+NC^{2}$ $K^{2}C+C^{2}$
Local Self Attention $\mathcal{O}(NK^{2}C+NC^{2})$ $5NC^{2}+NK^{2}C+N+2NK^{2}$ $5C^{2}$
Convolution $\mathcal{O}(NK^{2}C^{2}+NC^{2})$ $N2K^{2}C^{2}+NC^{2}$ $K^{2}C^{2}+C^{2}$
Global Self Attention $\mathcal{O}(N^{2}C+NC^{2})$ $5NC^{2}+N^{2}C+N+2N^{2}$ $5C^{2}$
4 Experiments
This section covers our experimental setup, describing the datasets and training routines. Since medical image datasets are often smaller than natural image datasets like ImageNet, we employ the smallest MetaFormer variant S12 with about 11.4 million trainable parameters (excluding the token mixers) to carry out our experiments. It comes with 64, 128, 320 and 512 channels for its four different stages. Each stage comprises two MetaFormer blocks, except for the third stage (six blocks, see Fig. 2(b)). For both settings, classification and segmentation, we compare the different token mixers introduced in Sec. 3.2 utilizing different pooling or kernel sizes where applicable (i.e. not for global attention and identity). Our code is available on GitHub222github.com/multimodallearning/MetaFormerMedImaging/tree/clean_code.
4.1 Classification
For classification, we extend the MetaFormer encoder with a classification head, averaging the final feature map to a vector via Global Average Pooling (GAP) and then linearly mapping it to the number of classes (see Fig. 2(b)). For optimization, we employ AdamW (Loshchilov and Hutter, 2019) with a weight decay of $0.1$ using the default learning rate of $1e^{-3}$, which is reduced by a single cycle of cosine annealing (Loshchilov and Hutter, 2017) (timm implementation) to a minimum of $1e^{-5}$. We found it beneficial to employ a linear learning rate warm-up during the first 5 epochs of training. We adapt the number of epochs to dataset size, targeting 35k iterations with batch size 128. The final model is selected based on the highest validation F1-score measured at the end of each training epoch. For the loss function, we chose cross-entropy with label smoothing of 0.1 and class weighting by the inverse root class frequency (clamped to a maximum of 10). Data augmentation consists of random affine transformations with a transformation matrix sampled from a zero-mean normal distribution with 0.1 standard deviation. For additional regularisation, we follow (Yu et al., 2022) utilizing stochastic depth (Huang et al., 2016) and LayerScale (Touvron et al., 2021). All images were bilinearly reshaped to $224\times 224$.
Given the scarcity of medical data, pretrained weights are often essential. This applies especially for self-attention, which converges slower and requires more data due to its weaker inductive bias compared to convolution (Lu et al., 2022). Consequently, we evaluate two MetaFormer variants: The first variant is trained from scratch, with the token mixer of all stages change to the specific experiment setting (denoted as $[T,T,T,T]$). Since there are no pretrained weights for all our deployed token mixers, we adapted the hybrid architecture of pooling, pooling, global self-attention, global self-attention, for which ImageNet weights have been made public (Yu et al., 2022), as our second pretrained variant. Here we keep the pooling in the first two stages and exchange the self-attention in the last two stages according to the experiment setting (denoted as $[P,P,T,T]$) and fine-tune the channel-MLP of all stages. Notably, the third stage contains six layers (vs. two in the others), so modifying it changes most of the token mixers. To measure the improvement provided by the pretrained weights, we also train the second variant ($[P,P,T,T]$) from scratch (result shown in appendix Tab. 6).
4.1.1 Datasets
We include various classification datasets with vastly different numbers of training images from multiple medical image modalities using the MedMNIST V2 collection Yang et al. (2023). From the collection, we select tasks that are harder to classify and thus suitable for comparing token mixers. We gauge the difficulty of the twelve MedMNIST2D datasets by training a ResNet18 with the aforementioned routine (Sec. 4.1) for up to 500 epochs or 35k35\,k iterations, without a learning rate scheduler. We then select datasets with a validation F1-score below 0.95 for our studies. This procedure yielded four diverse datasets: PathMNIST (Kather et al., 2019) provides patches from hematoxylin and eosin stained colorectal cancer histological images for the task of classifying nine different types of tissues. The images are rotation invariant and are split into 89 996/10 004/7 18089\,996/10\,004/7\,180 train/validation/test images, making it the largest dataset in our work. DermaMNIST (Tschandl et al., 2018; Codella et al., 2019) consists of dermatologic images of nine common pigmented skin lesions, leaving it also rotation invariant. The images are divided into 7 007/1 003/2 0057\,007/1\,003/2\,005 splits. With PneumoniaMNIST (Kermany et al., 2018), we include pneumonia detection in chest X-rays with 4 708/524/6244\,708/524/624 images, respectively. Lastly, in OrganSMNIST (Bilic et al., 2019; Xu et al., 2019) we classify eleven organs in the sagittal view of 201 CT scans, obtaining 13 932/2 452/8 82713\,932/2\,452/8\,827 images respectively.
To measure the factor of a potential domain shift between the pretrained ImageNet weights and the medical data, we also include a small subset of the ImageNet dataset, since a full-size ImageNet training would have been unfeasible considering the number of experiments and available hardware. We settle on ImageWoof (Howard, 2019): A non-trivial dataset of ten different dog breeds, coming with a predefined split of 9 035/3 9399\,035/3\,939 train/test images.
4.1.2 CNN Baseline
As a CNN baseline, we employ ResNet18 (He et al., 2016) (torchvision implementation), which matches the parameter count of MetaFormerS12. For models trained from scratch, we test multiple kernel sizes as in MetaFormer, while for pretrained models we use only kernel size 3, the only available pretrained variant.
4.2 Semantic Segmentation
For semantic segmentation, we extract the last block’s feature map of all four stages and employ the same light weighted all-MLP decoder used in the SegFormer (Xie et al., 2021) adding $525\,\mathrm{k}$ parameters to the MetaFormer. We chose this small decoder to minimise its impact on the segmentation tasks and have the MetaFormer be the main feature extractor, amplifying the potential performance difference between types of token mixers. Fig. 2(b) shows a schematic depiction of the decoder. The SegFormer decoder takes the feature maps from the four MetaFormer stages, equalizes the number of channels to 256 using linear layers, and upsamples them bilinearly to the first stage spatial resolution. A final two-layer channel-MLP with 256 hidden neurons then projects the concatenated $4\cdot 256$-channel feature map to the number of classes. Since, the feature map from the first stage is downsampled by factor four by the MetaFormer’s first patch embedding, the logits of the final predicted segmentation mask are bilinearly upsampled by factor four. It is worth noting, the MLP-decoder only includes pointwise operations and thus no token mixing takes place except within the bilinear upsampling (analogous to distance-weighted average pooling). We use the same training routine as for classification (see Sec. 4.1), only swapping the loss function to cross-entropy plus Dice loss (equally weighted) and setting a fixed number of epochs of 1000 regardless of the dataset size. Whenever a multiclass segmentation task includes the background class, we ignore it during the loss calculation.
4.2.1 Datasets
We use three different medical datasets for semantic segmentation. The Japanese Society of Radiological Technology (JSRT) Dataset (Shiraishi et al., 2000) is publicly available and consists of 247 chest X-rays. Each image comes with landmark annotations (Gaggion et al., 2022) for the four anatomical structures of the lungs, heart, and clavicles. For our experiments, we converted the landmarks to dense masks, downsample the dataset to $256\times 256$ pixels, and divide it into a custom split of 160/87 training/test images.
As a second dataset, we include bone segmentation of the paediatric wrist based on the publicly available GRAZPEDWRI-DX (GRAZ) dataset (Nagy et al., 2022). Here we used human expert segmentation mask of 17 bones provided in Keuth et al. (2025) of 63 images, which we downsample to $384\times 224$ and split into 43/20 for training/testing. As the last dataset, we include the Tumor InfiltratinG lymphocytes in breast cancER (TIGER) dataset [^3] TIGER training data was accessed on July 25th, 2025 from registry.opendata.aws/tiger., comprising large histology slices. Here, the goal is to segment seven different tissues, including different states/types of tumour and tumour-associated stroma. We selected the subset of slices where the region of interest is provided with dense segmentation mask, obtaining 284 images split into 256/28 training/testing. Since the slices have a very high resolution, we follow a patch-based approach with a patch size of $768\times 768$. During training, we sample one patch per image on-the-fly, with probabilities weighted by the inverse square root of class frequency to oversample minority classes. For inference, we apply a sliding window with $25\,\%$ overlap and merge predictions using Gaussian weighting, reducing the influence of uncertain patch edges.
4.2.2 CNN baseline
As a CNN baseline for semantic segmentation, we employ the U-Net (Ronneberger et al., 2015) with its extension of residual units (MONAI implementation). Each U-Net level employs the same number of channels as the corresponding MetaFormerS12 stage (64, 128, 320, 512). To allow the U-Net to have a sufficiently large receptive field, we deploy an additional fifth stage with 1024 channels. We construct U-Nets with $K\in\{3,5,7\}$ and additionally $K=9$ for the TIGER dataset because of its larger image size. We match the U-Nets’ number of parameters to be as close as possible to the corresponding MetaFormer with convolutional token mixers by varying the number of residual units at each stage. To make sure, each stage is capable of extracting features, we set a minimum of one residual unit.
Since the MetaFormer’s first patch embedding uses a stride of four, its predicted segmentation map is coarser and must be upsampled accordingly (Sec. 4.2). For better comparison, we introduce a second U-Net variant denoted as UNet@PatchEmb applied to a four times downsampled image by the same layer as the MetaFormer’s first patch embedding. This version omits the fifth U-Net stage to match the MetaFormer’s receptive field and adjusts the number of residual units to obtain a comparable parameter count. Like in the MetaFormer, the predicted segmentation mask is upsampled by a factor of four to match the input resolution.
4.3 Ranking
To determine the overall best performing token mixer in each setting, we implement a ranking scheme based on the well-established challenges Medical Decathlon (Antonelli et al., 2022) and Learn2Reg (Hering et al., 2022). Intuitively, every token mixer with a specific kernel size can be considered an individual challenge submission. On each dataset, token mixers are compared against each other and earn a point, for each comparison they significantly won. The results are sorted and token mixers are given scores from the range $[0.1,1]$ per dataset. In cases of ties, they share the average rank score. Finally, an aggregated global rank is computed by the geometric mean across datasets. Since for classification tasks, there is no instance-level performance with the Area under the Receiver Operating Characteristic curve (AUC) metric, we used bootstrapping of the test data with 5k5\,k repeats. With it, we estimate the two-sided $95\,\%$ confidence interval of the difference in AUC to determine the significance of a token mixers’ comparison. For segmentation, we utilize the Wilcoxon signed-rank test with $p<0.05$ on Dice Similarity Score (DSC).
5 Results and Discussion
5.1 Classification
Table 2: Area under the Receiver Operating Characteristic curve for different token mixers $T$ across datasets. The first meta row holds our models trained from scratch ($[T,T,T,T]$). The second utilizes ImageNet-pretrained weights for the channel-MLPs and only exchanges token mixers in the last two stages ($[P,P,T,T]$). Number of trainable parameters (#P) is given in millions. See Sec. 4.3 for ranking schema details. Refer to Fig. 4(b) for the impact of pre-training.
* = reduced learning rate and doubled epochs; + = $\frac{\text{batch size}}{4}$ but gradient accumulation of 4; $\dagger$ disabled compilation (cf. Sec. 5.1.4).
Dataset ImageWoof Path Derma Pneumonia OrganS
(# training images) ($9\,\mathrm{k}$) ($90\,\mathrm{k}$) ($7\,\mathrm{k}$) ($6\,\mathrm{k}$) ($14\,\mathrm{k}$)
Token Mixer $T$ Kernel MedMNIST Rank #P [M]
$[T,T,T,T]$ from scratch
pooling 3 0.9495 0.9758 0.8883 0.9921 0.9617 0.691 11.4
5 0.9517 0.9799 0.8911 0.9832 0.9375 0.484 11.4
7 0.9539 0.9829 0.8795 0.9814 0.9420 0.563 11.4
conv 3 0.9483 0.9834 0.8675 0.9761 0.9534 0.541 22.0
5 0.9353 0.9878 0.8855 0.9807 0.9475 0.616 40.9
7 0.9235 0.9887 0.8628 0.9925 0.9503 0.563 69.2
grouped conv 3 0.9534 0.9803 0.8867 0.9886 0.9535 0.767 11.4
5 0.9412 0.9821 0.8746 0.9863 0.9410 0.508 11.5
7 0.9341 0.9798 0.8681 0.9878 0.9440 0.477 11.6
$\text{local attn}^{*}$ 3 0.9152 $0.9701^{\dagger}$ 0.8869 $0.9739^{\dagger}$ 0.9384 0.34 16.1
5 0.9104 $0.9726^{\dagger}$ 0.8538 0.9780 $0.9274^{\dagger}$ 0.205 16.1
7 0.9014 $0.9782^{\dagger}$ 0.8458 0.9803 $0.9326^{\dagger}$ 0.259 16.1
$\text{global attn}^{+}$ - 0.8885 0.9802 0.9174 0.9493 0.9679 0.353 16.5
identity 1 0.9575 0.9756 0.8784 0.9883 0.9626 0.609 11.4
$[P,P,T,T]$ pretrained channel-MLPs
pooling $P$ 3 0.9645 0.9596 0.9064 0.9830 0.9536 0.366 11.4
5 0.9641 0.9651 0.9191 0.9883 0.9618 0.474 11.4
7 0.9697 0.9750 0.9198 0.9793 0.9475 0.477 11.4
conv 3 0.9680 0.9564 0.9127 0.9832 0.9545 0.401 21.7
5 0.9563 0.9678 0.9132 0.9861 0.9609 0.405 39.9
7 0.9607 0.9721 0.9143 0.9880 0.9567 0.409 67.2
grouped conv 3 0.9642 0.9620 0.9000 0.9906 0.9440 0.422 11.4
5 0.9556 0.9641 0.9046 0.9918 0.9545 0.396 11.5
7 0.9831 0.9800 0.9014 0.9864 0.9438 0.514 11.6
local attn 3 0.9671 0.9745 0.9097 0.9801 0.9489 0.462 16.0
5 0.9654 0.9742 0.9043 0.9935 0.9568 0.595 16.0
7 0.9666 0.9736 0.9197 0.9755 0.9554 0.533 16.0
local attn (warm start) 3 0.9723 0.9753 0.9030 0.9814 0.9582 0.613 16.0
5 0.9904 0.9729 0.9032 0.9913 0.9449 0.565 16.0
7 0.9919 0.9796 0.9028 0.9894 0.9624 0.695 16.0
global attn - 0.9646 0.9799 0.9046 0.9887 0.9512 0.546 16.0
(warm start) - 0.9920 0.9847 0.9067 0.9915 0.9314 0.526 16.0
identity 1 0.9840 0.9511 0.9646 0.9911 0.9574 0.543 11.4
ResNet18
from scratch 3 0.9381 0.9735 0.9135 0.8885 0.9627 11.2
5 0.9355 0.9735 0.8772 0.9833 0.9558 30.7
7 0.9165 0.9772 0.8639 0.9868 0.9378 60.0
pretrained 3 0.9875 0.9871 0.9476 0.9390 0.9558 11.2
Figure 3: Area under the Receiver Operating Characteristic curve for MetaFormer trained from scratch with different token mixers (architecture signature $[T,T,T,T]$). Line plots indicate results with a monotonic change across kernel sizes.
5.1.1 Dataset-specific Performance of Token Mixers
Fig. 3 and Tab. 2 (meta row $[T,T,T,T]$) show the Area under the Receiver Operating Characteristic curve (AUC) of MetaFormers trained from scratch utilizing the specified token mixer at each four stages.
For ImageWoof, token mixers with additional learnable parameters decrease the performance, instead identity and pooling yield the best AUC.
On PathMNIST, which consists of pathology slide images where a larger spatial context is advantageous, we observe a positive effect from larger kernel size on the AUC, while the identity mapping yields lower performance. Since PathMNIST provides $90\,\mathrm{k}$ training images, it allows for the effective optimization of models with many parameters. This is reflected in the performance of the convolutional token mixers, which achieve the overall best results with a kernel size of 7. This configuration adds the highest amount of trainable parameters with a large margin ($57.8\,\mathrm{M}$ parameters on top of the base $11.4\,\mathrm{M}$).
On DermaMNIST, global self-attention shows by far the best performance. For the other token mixers, we observe a pattern similar to ImageWoof, where larger kernels and increased parameter counts appear to hinder generalization. Because pneumonia-infected tissue can span the entire lung in chest X-rays of PneumoniaMNIST, a larger receptive field should be beneficial. The result for convolution is consistent with this hypothesis. However, for pooling and grouped convolution, different kernel sizes yield comparable AUC. Moreover, global attention with the largest possible receptive field performs worst among all token mixers in this task. Nevertheless, as all token mixers (except global attention) achieve an AUC close to the optimum of 1.0, pneumonia detection can be regarded as a relatively easy classification task. Hence, all models successfully learn discriminative features and, thus, the effect of an increased capacity (more parameters or larger kernels) is less noticeable.
For OrganSMNIST, which involves organ classification in abdominal CT slices, no clear trend emerges regarding neither the optimal kernel size nor choice of token mixing. The highest AUC is reached by global attention, followed pooling with $K=3$.
When comparing the correlation between kernel size and performance across datasets, one might observe a trend: For datasets such as ImageWoof and DermaMNIST, which typically contain a single, localized object per image, token mixing within the immediate neighbourhood tends to be more effective. Conversely, when the classification task requires capturing broader spatial context as in PathMNIST (pathology slides) or PneumoniaMNIST (chest radiographs, where infected tissue can span the entire lung) – token mixers with a larger receptive field provide some advantage.
5.1.2 Global Ranking of Token Mixers
Grouped convolution with a kernel size of $K=3$ obtains the overall best rank (0.767), followed by pooling with $K=3$ (0.691). To make a more general statement, we averaged the rankings over the different token mixers and kernel sizes (see right column in Tab. 2). We observe that less complex token mixers perform better on average, putting grouped convolution (0.584) first, followed by pooling (0.579), convolution (0.573) and with a large margin local attention (0.268). Moreover, smaller kernel sizes appear to be favourable, with $K=3$ (0.585) and 5 and 7 comparable (0.453/0.466). So in general, less complex token mixers with a smaller kernel sizes perform the best on average across different data sets. See appendix Tab. 7 for absolute and relative ranking scores for each dataset. The noticeably lower ranks of attention-based token mixers with local on average of 0.268 and global with 0.353 suggest them less suitable when training from scratch on relatively small datasets. However, due to local attention’s training instabilities (see Sec. 5.1.4), we had to reduce the learning rate to $1e^{-4}$. Although, we doubled the number of epochs for a comparable overall optimization budget, the reported AUCs could still be underestimating local attention’s potential. The better average rank of grouped convolution (0.584) than convolution (0.573) suggests that the channel cross-talk lacked by grouped convolution can be effectively compensated by the channel-MLP within the MetaFormer block. Consequently, grouped convolution should be preferred, as it offers lower runtime complexity and reduced parameter count regarding the number of channels (cf. Tab. 1). No token mixing (identity) reaches a competitive place of 4/14 (rank 0.609), which is consistent with the findings of Yu et al. (2023). This raises whether additional token mixing is necessary at all, as the convolution (without any non-linearity) used for patch embedding at the beginning of each stage may already provide sufficient token mixing. Another contributing factor could be the MetaFormer’s extension for global prediction: the final feature map is collapsed via Global Average Pooling (GAP) into a single vector (cf. Fig. 2(b)), which may itself introduce enough token mixing for classification. Furthermore, because each MetaFormer block includes skip connections (cf. Fig. 1), the token mixing by GAP may propagate transitively through all stages of the network.
5.1.3 Transfer Learning from ImageNet
Another factor that may explain the poor performance of local self-attention, which ranked last in the previous experiment, is its slow convergence and high data requirements as reported in Lu et al. (2022). To mitigate this limitation, we experiment with initializing our models with publicly available ImageNet-pretrained weights. However, since pretrained weights are not available for all our deployed token mixers, we adopt a hybrid MetaFormer variant that uses pooling in the first two stages and global attention in the last two (Yu et al., 2022). To avoid a drastic domain shift, we keep pooling as token mixer of the first two stages (introducing no token mixer-parameters) and replace only the global attention in the last two stages with our token mixers. This architecture signature is denoted as $[P,P,T,T]$ (see the corresponding meta row in Tab. 2). Because replacing the token mixer does not modify the channel-MLP, we can leverage pretrained weights for all stages. For self-attention (global and local), we differentiate between two variants: First, transfer the pretrained attention weight matrices $\mathbf{W}_{\{k,v,q,u\}}$ (cf. Sec. 3.2.5) calling it “warm start” and second, initializing their parameters randomly. Since the architecture signatures between the two experiments differs ($[T,T,T,T]$ vs. $[P,P,T,T]$), we also train a $[P,P,T,T]$ variant from scratch (results shown in appendix Tab. 6), allowing us to measure the impact of pretrained weights.
Fig. 4(a) shows the absolute performance in AUC of the $[P,P,T,T]$ architecture utilizing pretrained weights, and Fig. 4(b) shows the relative improvement when compared to its counterpart trained from scratch. The performance of warm start-local attention on ImageWoof demonstrates that the attention weights of global attention can be effectively reused. As expected, this effect becomes more pronounced as the kernel size of local attention increases, making the two operations more similar. Beyond the benefit of reusing attention weights ($10.54\,\%$, $13.95\,\%$, $15.53\,\%$), employing pretrained weights for the channel-MLPs alone also yields a substantial improvement ($9.94\,\%$, $11.08\,\%$, $12.13\,\%$) for local attention trained from scratch. The performance gain is as strong for global attention with ($11.37\,\%$) and without warm start ($8.29\,\%$). We also observe performance gains from pre-training across all other token mixers on ImageWoof even if the introduced domain shift diminishes this effect. Notably, the magnitude of improvement correlates with the similarity to the original global self-attention token mixer: for example, parameterized convolutions with the largest kernels possibly benefit more than non-parameterized pooling or no token mixing (identity).
In the medical setting, the impact of ImageNet pretrained weights is far less pronounced. On PathMNIST, we observe the same positive correlation for larger context as we have in the $[T,T,T,T]$ variant trained from scratch. However, the relative benefit of using pretrained weights is limited and is only noticeable for attention-based token mixers with kernel size $K\geq 5$. On DermaMNIST, no toking mixing (identity) outperforms any token mixing operation by a large margin (AUC of 0.96, second best are pooling and local attention with $K=7$ at 0.92). Beyond its strong absolute performance, identity mapping does not benefit much from pretrained weights on DermaMNIST. Attention-based mixers even perform worse when pretrained compared to training from scratch on this dataset. In contrast, convolution- and pooling-based token mixers gain the most, with larger kernels showing greater improvements. These improvements are the largest among the medical datasets. We attribute this to the dataset’s domain – colour microscope images of skin lesions, typically a single localised object – which is closer to the ImageNet domain and suits pooling and convolutional layers. Additionally, because DermaMNIST has the lowest AUCs compared to the other datasets, it offers the greatest potential for performance improvement. In PneumoniaMNIST, we find the same positive correlation between kernel size and absolute performance and a benefit of up to $2.5\%$ improvement when using pretrained weights for all token mixers except pooling. As in the previous experiment, we do not observe any clear performance dependencies on the OrganSMNIST dataset. Here, most token mixers yield worse performance when using pretrained ImageNet weights compared to training from scratch. When averaging the relative improvements of all token mixers across the medical datasets, convolution achieves the largest gain ($0.9\,\%$), while global attention with warm-start on average does not benefit from pretrained weights ($0.8\,\%$). These mixed results imply that ImageNet pretrained weights can only be applied to medical image classification with some caveat. However, the mean improvement per kernel size $K$ further supports our interpretation, that moving closer to the receptive field of the original token mixer (global attention) helps narrow this domain gap, with $K=7$ yielding $0.59\,\%$, $0.56\,\%$ for $K=5$, and $0.26\,\%$ with $K=3$. Furthermore, the average improvement of $7.06\,\%$ on ImageWoof indicates that transferring pretrained weights to different token mixers is very effective when staying within the natural image domain.
When considering the global ranks, we find 7/8 of the first ranks (18 entries in total, see Tab. 2) holding attention-based token mixers. Local attention with warm start initialization with kernel size $K=7$ and $K=3$ hold the first (0.695) and second (0.613) place, respectively. We attribute the lower ranks of non-attention-based token mixers to the domain gap introduced by replacing the original attention operation, hindering the benefit from pretrained channel-MLPs. However, the relatively low rank (0.526) of global attention with warm-start initialization (8th place) suggests that the superior performance of local attention cannot be explained solely by its similarity to the pretrained global attention, but also by the presence of a local inductive bias. For the non-attention-based token mixers, we observe the same pattern as before, ranking less complex token mixer higher. Grouped convolution and pooling rank similarly, with an average rank of 0.444 and 0.439, followed by convolution (0.405). Interestingly, identity achieves a competitive rank of 0.5426, making it the only non-attention-based token mixer in the top eight. This suggests again that token mixing may be of secondary importance in a global prediction setting, where the final feature map is spatially averaged into a single vector before the classification head. See appendix Tab. 7 for absolute and relative ranking scores for each dataset.
((a)) Area under the Receiver Operating Characteristic curve for the MetaFormer using pretrained ImageNet weights utilizing different token mixers in the last two stages.
((b)) Relative improvement from utilizing pretrained weights compared to the variant trained from scratch.
Figure 4: Performance of the MetaFormer using pretrained ImageNet weights for its channel-MLPs and different token mixers in the last two stages (architecture signature $[P,P,T,T]$). Line plots indicate results with a monotonic change across kernel sizes.
5.1.4 Training Instabilities of Flex Attention
We employ PyTorch’s flexattention framework to implement the necessary attention pattern for local attention (cf. Eq. (8)). The framework heavily relies on torch.compile to reduce both memory consumption and computational overhead. By monitoring the per-layer gradient $L_{2}$-norm, we observe that compilation can induce gradient explosions within the flexattention module when training on large input resolutions. For classification, this limit is at the spatial resolution of MetaFormer’s first stage (in our setting $56\times 56$). To stabilize training, we experimented with gradient clipping, returning to full 32 bit precision (we usually utilize 16 bit bfloat for efficiency) and also prescaling $\mathbf{q}$ by $\frac{1}{\sqrt{D}}$ before its correlation with $\mathbf{k}$ to reduce the magnitude of values within the matrix multiplication (cf. Eq. 8). In the end, reducing the learning rate from $1e^{-3}$ to $1e^{-4}$ solved the issue for the most situations. To keep the training budget comparable, we double the number of epochs (marked with * in Tab. 2). In cases, where reducing the learning rate was not sufficient, we are forced to disable the compilation of flexattention (marked with $\dagger$ in Tab. 2). This massively increases the resource consumption from one hour training time with around $10\,\text{GB}$ VRAM at a batch size of 128 to 20 hours and $60\,\text{GB}$ at a batch size of 64. To compensate for the reduced batch size, we employ gradient accumulation to simulate an update with the original batch size. Interestingly, we do not encounter any issues when employing the MetaFormer as encoder for segmentation regardless of the larger input sizes up to $768\times 768$ for the TIGER task. We attribute this stability to the SegFormer decoder, which adds shortcuts to each MetaFormer stage (cf. Fig. 2(b)) and thereby shortens the effective gradient path.
5.1.5 ResNet Baseline
The ResNet18 baseline trained from scratch gets outperformed by its MetaFormer counterpart (employing convolutional token mixer in $[T,T,T,T]$) on PathMNIST (by $1.2\,\%$ AUC) and PneumoniaMNIST (by $3\,\%$ AUC) on average. The same holds for grouped convolution with $0.5\,\%$ and $3.5\,\%$ AUC, respectively. On the two more challenging medical datasets (DermaMNIST and OrganSMNIST), the ResNet18 performs on par with the MetaFormer. Overall, we observe the ResNet18 following the same trends for different kernel sizes across datasets as we saw in the MetaFormer.
For the ResNet18 employing ImageNet pretrained weights, we find a similar pattern implying that ImageNet pretrained weights do not provide a substantial improvement on MedMNIST: using pretrained ImageNet weights yields the largest gains on ImageWoof ($5.27\,\%$), DermaMNIST ($3.73\,\%$), and PathMNIST ($1.01\,\%$), but results in a performance drop on PneumoniaMNIST ($4.84\,\%$) and OrganSMNIST ($0.69\,\%$).
5.2 Segmentation
Table 3: Dice Similarity Score (DSC) across three medical datasets for semantic segmentation. We exchange the token mixers at all four stages. All models are trained from scratch. See Sec. 4.3 for ranking schema details. Number of trainable parameters (#P) is given in millions.
Dataset (number of training images) JSRT (160) GRAZ (43) Tiger@768² (256)
TokenMixer $T$ Kernel Rank # P[M]
$[T,T,T,T]$ from scratch
pooling 3 0.9468 0.8094 0.6125 0.608 11.9
5 0.9463 0.8072 0.5555 0.478 11.9
7 0.9450 0.7973 0.5580 0.406 11.9
9 0.6019 11.9
conv 3 0.9486 0.8164 0.6024 0.793 22.6
5 0.9510 0.8276 0.5622 0.687 41.4
7 0.9504 0.8330 0.5623 0.707 69.7
9 0.5689 107
grouped conv 3 0.9500 0.8119 0.5791 0.596 12.0
5 0.9495 0.8319 0.5941 0.811 12.0
7 0.9504 0.8320 0.5711 0.832 12.1
9 0.5893 12.2
local attn 3 0.9418 0.8165 0.5167 0.212 16.7
5 0.9465 0.7957 0.5303 0.374 16.7
7 0.9449 0.7927 0.5374 0.296 16.7
9 0.5511 16.7
global attn - 0.9443 0.7562 DNS 0.192 16.7
identity 1 0.9458 0.7417 0.5358 0.198 11.9
CNN baseline
UNet 3 0.9552 0.8478 0.5666 28.4
5 0.9499 0.8258 0.5458 42.1
7 0.9505 0.8318 0.5533 82.1
9 0.5544 135
UNet@PatchEmb 3 0.9357 0.7738 0.5735 21.2
5 0.9342 0.7864 0.5981 39.4
7 0.9335 0.7629 0.6047 77.1
9 0.5639 96.6
Figure 5: Dice Similarity Score across three medical datasets for semantic segmentation, investigating into different token mixer and kernel sizes. Line plots indicate performance with a monotonic change across kernel sizes.
5.2.1 Dataset-specific Performance of Token Mixers
For semantic segmentation as a dense prediction task, we employ the MetaFormer as image encoder with different token mixers at every stage, hence having an architecture signature of $[T,T,T,T]$. We extend its architecture with a lightweight MLP-only decoder of the SegFormer (Xie et al., 2021) (see Fig. 2(b) and Sec. 4.2 for details). Tab. 3 and Fig. 5 show the average Dice Similarity Score (DSC) across three datasets for different token mixers.
Segmentation of the lungs, heart, and clavicles in JSRT chest X-rays constitutes a relatively straightforward task, as all evaluated models achieve a Dice Similarity Score (DSC) of at least 0.94. As a result, this dataset does not provide informative distinctions in performance between different token mixers or kernel sizes.
Next, we consider the GRAZ dataset, which involves segmenting 17 bones in paediatric wrist radiographs and presents challenges such as small, overlapping carpal bones. The results highlight the importance of an inductive bias with strong locality, as convolutional token mixers with the kernel size $K>3$ outperform pooling- and attention-based alternatives. Here, larger kernels further improve performance, with $K=7$ yielding an increase of nearly 2 DSC percent points compared to $K=3$ ($83.3\,\%$ vs. $81.6\,\%$). Global self-attention performs worst, indicating that while a large receptive field is available, high local inductive bias is critical for this task. However, the overall performance of attention-based token mixers may be constrained by the limited dataset size: 43 training images may be insufficient to reliably learn attention patterns.
We use the opportunity provided by the large histology slides of the TIGER dataset to employ an additional kernel size $K=9$ to evaluate if and to which extent a larger receptive field is beneficial for such tasks. As it is common for histology tasks with high-resolution images, we settle for a patch-based approach using patches of size $768\times 768$ (see Sec. 4.2 for details). While we observe a positive correlation between local self-attention performance and a larger receptive field size, no such clear trend exists for the other token mixers. It would have been informative to evaluate whether this trend extends to the global receptive field by global attention; however, the sequence length resulting from our patch size ($768^{2}=589\,824$) exceeds the capacity of even our largest available GPU (NVIDIA H100 with $94\,\text{GB}$ VRAM) regardless of employing gradient accumulation reducing the batch size down to 1 and employing gradient checkpointing after each MetaFormer block. Although no consistent pattern emerges for the other token mixers, all of them outperform even the best-performing local attention with $K=9$.
5.2.2 Global Ranking of Token Mixers
The global rank scores in Tab. 5 (right column) suggest that dense prediction tasks require explicit token mixing at every stage, which was not the case for global prediction tasks (cf. Sec. 5.1.2). We find grouped convolution with a large kernel size $K=7$ yielding the highest global rank (0.832) followed by its $K=5$ variant (0.811). However, we cannot observe the pattern of larger kernels yielding better ranks for the other token mixers. Furthermore, considering the locality of the ranked token mixers (averaged global rank over kernel sizes), from the high locality of grouped/standard convolution (0.746/0.729) to pooling (0.498), local attention (0.294), and finally global attention (0.192), it is evident that a strong local inductive bias is crucial for segmentation in addition to the token mixing itself. This lack of inductive bias hinders the performance of global attention, especially for the small datasets at hand. The identity operation ranks similarly low as global attention does (0.198). As already observed in the classification experiments, using standard convolution (rank 0.729) as the token mixer introduces unnecessary runtime and parameters, since the channel cross-talk that grouped convolution (rank 0.746) would lack is already provided by the MetaFormer block’s channel-MLP. See appendix Tab. 8 for absolute and relative ranking scores for each dataset.
5.2.3 U-Net Baselines
For the CNN baseline results, see Tab. 3. The U-Net exhibits a negative correlation between performance and kernel size, suggesting overfitting due to high number of parameters at larger kernels on our limited medical datasets. However, the variant with kernel size $K=3$ outperforms the best-performing MetaFormer on the JSRT and GRAZ datasets. Since the initial patch embedding in the MetaFormer downsamples the input by a factor of four, the predicted segmentation mask is coarser compared to the U-Net. For a more fair comparison, we also evaluate a U-Net applied to the same downsampled patch embedding (UNet@PatchEmb, see Sec. 4.2). In this setting, U-Net performs noticeably worse than the MetaFormer with convolutional token mixers. While we observe a positive correlation with larger kernels on the TIGER dataset for UNet@PatchEmb with kernel size $K\leq 7$, the high number of parameters ($96.6\,\text{M}$) associated with $K=9$ may be difficult to train effectively given the limited dataset size. Note that for each of the 256 TIGER train images, multiple $(768\times 768)$-sized patches can be extracted, increasing the effective dataset size.
6 Conclusion and Future Work
In this work, we systematically compared different token-mixing operations – pooling, (grouped) convolution, local and global self-attention – within the MetaFormer architecture across eight datasets, seven from the medical imaging and one from the natural image domain. We evaluated two task categories: global prediction, using five image classification datasets, and dense prediction, using three semantic segmentation datasets, comparing our results against CNN-based baselines.
When comparing different kernel sizes for classification, we observe that datasets containing a single dominant object in the image (e.g., ImageWoof and DermaMNIST) favour more locality and thus benefit from smaller kernel sizes. In contrast, tasks with larger regions of interest, such as PathMNIST and PneumoniaMNIST, tend to profit from incorporating a broader spatial context through larger kernels. Based on our ranking schema, we found simpler token mixers like grouped convolution and pooling favourable when training from scratch. Besides their good performance, their low runtime complexity and small amount of trainable parameter (if any) make them a sensible choice. Additionally, even omitting token mixing entirely (using the identity operation) yields comparable classification performance. We attribute this to MetaFormer’s classification extension, which employs Global Average Pooling (GAP) on the final feature map, effectively providing a sufficient implicit token-mixing. For in-domain evaluation on ImageWoof, we further showed that transferring ImageNet-pretrained channel-MLPs is beneficial, even when replacing the MetaFormer token mixer. Our analysis reveals, however, that a token mixer that is more similar to the original one (global attention) will benefit more from pretraining, since a change of token mixer introduces a domain shift. Specifically, local self-attention utilizing large kernels benefit most, as its receptive fields approaches a global context similar to global attention. We additionally demonstrate that transferring the attention weights from global to local attention provides further gains. This proofs local self attention as an attractive alternative for global self attention in a setting with pretrained weights available. Especially when dealing with large input resolutions, where its local receptive provides a lower runtime complexity. However, when transitioning to the medical domain, we find that the domain gap is typically too large for ImageNet-pretrained weights to offer a consistent improvement across datasets in our setting. Therefore, releasing pretrained models trained on a diverse set of medical datasets could provide a valuable contribution to the community in the future.
For segmentation tasks, we observe that a strong local inductive bias is crucial, making convolutional token mixers the most effective choice. Here, grouped convolution achieves the highest rank across datasets while offering lower runtime complexity and fewer parameters than standard convolution. We attribute this superior performance to the channel-MLPs, which already provide the channel cross-talk that gets omitted in grouped convolution. Because our segmentation decoder (a lightweight SegFormer design (Xie et al., 2021)) does not introduce token mixing, this must be provided within the MetaFormer. Consequently, using no token mixing (identity) leads to inferior performance (second to last global rank). We explicitly chose a lightweight decoder to emphasize the impact of different token mixers. Future work might explore alternative decoder designs that incorporate varying degrees of token mixing like MetaSeg Kang et al. (2024).
To address the central research question of this work – how to mix tokens (Shaken or Stirred?) within the MetaFormer for medical image processing – our findings suggest that low-complexity token mixers such as grouped convolution, pooling, or even the identity operation are sufficient for classification tasks. We observe no major performance gains from using more complex mixers when training from scratch, confirming that the insights from MetaFormer studies on natural images (Yu et al., 2023, 2022) also generalize to the medical imaging domain. For segmentation tasks, however, we find strong locality crucial and therefore recommend grouped convolution as the preferred token mixer.
Since the medical domain also encompasses volumetric data (e.g., CT or MRI), extending the MetaFormer architecture to 3D inputs represents an important direction for future work. In 3D, the cubically scaling complexity of some token mixers may lead to other recommendations. While our study has focused on classification and segmentation as representative global and dense prediction tasks, other applications such as object detection are also of importance in the medical domain. Future work could explore using the MetaFormer as an image encoder within the Detection Transformer (DETR) framework (Carion et al., 2020).
[…]
This research has been funded by the state of Schleswig-Holstein, Germany, Grant Number 22023005.
[…]
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
[…]
We declare we don’t have conflicts of interest.
[…]
All the data used in this work are publicly available and can be accessed through the cited sources. To reproduce the preprocessing steps and our results, please follow the instructions provided in our GitHub repository: https://github.com/multimodallearning/MetaFormerMedImaging/tree/clean_code.
Declaration on Generative AI
During the preparation of this work, the authors used ChatGPT to paraphrase and reword. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the publication’s content.
Appendix A Overview of Appendix
To provide additional insights into our results, we include both accuracy and F1-score as supplementary metrics for the classification tasks. The F1-score might offer more intuition than the AUC, as it represents the harmonic mean of precision and recall, making it a more interpretable indicator of classification performance. Nevertheless, since the AUC evaluates classifier performance across various confidence thresholds, we consider it the primary metric for discussing the classification benchmarks in our work. In addition, we report the absolute and relative ranking scores of all token mixers across datasets for classification and segmentation. The following section provides an overview of the included tables and highlights their aspects that may be of particular interest to the reader.
  • Tab. 4 extends the evaluation of classifiers trained from scratch (MetaFormer with architecture signature $[T$, $T$, $T$, $T]$ and ResNet18) by including accuracy and F1-score. This table provides additional insights into the results discussed in Sec. 5.1.1.
  • Tab. 5 extends the evaluation of classifiers utilizing ImageNet-pretrained weights (MetaFormer with architecture signature $[P,P,T,T]$ and ResNet18) to include accuracy and F1-score. It further complements the discussion presented in Sec. 5.1.3.
  • Tab. 6 reports the AUC and F1-score for the MetaFormer $[P,P,T,T]$ variant trained from scratch. Based on these results and those from Tab. 2 (meta-row $[P,P,T,T]$), the percentage uplift values when compared to pretrained weights are calculated as shown in Fig. 4(b).
  • Tab. 7 presents the absolute and relative ranking scores of different token mixers across classification datasets. This information may be particularly useful for readers seeking the best-performing token mixers for their specific domains.
  • Similar to Tab. 7, Tab. 8 reports the absolute and relative ranking scores of token mixers across segmentation datasets.
Table 4: Performance for different token mixers $T$ across datasets. Each stage of the MetaFormer utilizes the given token mixer and was trained from scratch. Due to training instabilities within the implementation of local attention, we had to adapt the training routine (cf. Sec. 5.1.4). Please refer to markers ($\dagger$) in Tab. 2 metarow $[T,T,T,T]$ to identify those results.
ImageWoof PathMNIST DermaMNIST PneumoniaMNIST OrganSMNIST
Token Mixer $T$ Kernel Size Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
pooling 3 0.7908 0.7907 0.8817 0.8767 0.6828 0.6868 0.9632 0.9656 0.7659 0.7689
5 0.7941 0.7965 0.8832 0.8773 0.6933 0.6855 0.9393 0.9446 0.7494 0.7576
7 0.8044 0.8059 0.9030 0.8986 0.6811 0.6991 0.9526 0.9537 0.7649 0.7706
conv 3 0.7869 0.7891 0.8862 0.8814 0.7083 0.7123 0.9064 0.9201 0.7721 0.7765
5 0.7540 0.7567 0.9045 0.9035 0.7012 0.7013 0.9419 0.9480 0.7695 0.7743
7 0.7078 0.7090 0.9115 0.9112 0.6883 0.6872 0.9500 0.9535 0.7872 0.7901
grouped conv 3 0.7915 0.7937 0.8888 0.8840 0.7138 0.7029 0.9474 0.9532 0.7837 0.7863
5 0.7694 0.7713 0.9068 0.9065 0.7091 0.7062 0.9534 0.9569 0.7961 0.8001
7 0.7411 0.7424 0.8927 0.8884 0.6882 0.7121 0.9521 0.9552 0.7850 0.7871
local attn 3 0.7357 0.7371 0.8580 0.8476 0.7255 0.7263 0.9521 0.9521 0.7852 0.7868
5 0.7203 0.7206 0.8868 0.8781 0.6999 0.7199 0.9333 0.9408 0.7830 0.7875
7 0.6914 0.6908 0.8794 0.8728 0.6920 0.7081 0.9346 0.9395 0.7826 0.7866
global attn - 0.5439 0.5408 0.8778 0.8647 0.6046 0.5384 0.8688 0.8810 0.7357 0.7410
identity 1 0.7695 0.7698 0.8913 0.8822 0.6722 0.6761 0.9432 0.9466 0.7687 0.7720
ResNet18 3 0.7454 0.7478 0.8392 0.8260 0.7332 0.7388 0.8885 0.9048 0.7866 0.7904
5 0.7257 0.7271 0.8732 0.8615 0.7002 0.7331 0.9308 0.9404 0.7836 0.7907
7 0.6909 0.6928 0.8619 0.8496 0.6912 0.7015 0.9124 0.9241 0.7763 0.7832
Table 5: Performance for different token mixers $T$ across datasets. The MetaFormer utilizes ImageNet-pretrained weights for the channel-MLPs and only exchanges token mixers in the last two stages, keeping the original pooling at the first two stages ($[P,P,T,T]$).
ImageWoof PathMNIST DermaMNIST PneumoniaMNIST OrganSMNIST
Token Mixer $T$ Kernel Size Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
pooling $P$ 3 0.8501 0.8531 0.8643 0.8539 0.7585 0.7737 0.9321 0.9421 0.7620 0.7657
5 0.8610 0.8637 0.8726 0.8647 0.7232 0.7625 0.9521 0.9584 0.7632 0.7648
7 0.8612 0.8633 0.8914 0.8869 0.7789 0.8058 0.9423 0.9496 0.7725 0.7750
conv 3 0.8494 0.8524 0.8768 0.8687 0.7523 0.7624 0.9444 0.9513 0.7889 0.7908
5 0.8361 0.8383 0.8818 0.8705 0.7642 0.7894 0.9372 0.9459 0.7963 0.7967
7 0.8259 0.8256 0.8684 0.8570 0.7863 0.7818 0.9517 0.9568 0.7740 0.7794
grouped conv 3 0.8491 0.8521 0.8601 0.8475 0.7740 0.7936 0.9487 0.9549 0.7818 0.7831
5 0.8308 0.8338 0.8977 0.8949 0.7594 0.7928 0.9598 0.9654 0.7916 0.7973
7 0.8375 0.8400 0.8967 0.8919 0.7605 0.7916 0.9214 0.9330 0.7910 0.7965
local attn 3 0.8735 0.8753 0.8856 0.8775 0.7737 0.7923 0.9321 0.9421 0.7747 0.7790
5 0.8614 0.8619 0.8852 0.8786 0.7631 0.7613 0.9513 0.9552 0.7728 0.7790
7 0.8679 0.8697 0.8868 0.8784 0.7577 0.7738 0.9513 0.9583 0.7617 0.7667
local attn (warm start) 3 0.8742 0.8759 0.8918 0.8876 0.7899 0.7995 0.9607 0.9654 0.7769 0.7814
5 0.8841 0.8875 0.8884 0.8829 0.7879 0.8035 0.9632 0.9688 0.7921 0.7956
7 0.8834 0.8841 0.8991 0.8931 0.7778 0.7895 0.9436 0.9482 0.7862 0.7918
global attn - 0.8641 0.8662 0.8995 0.8962 0.7451 0.7714 0.9286 0.9386 0.7667 0.7732
global attn (warm start) - 0.8908 0.8939 0.9011 0.8990 0.7672 0.7821 0.9444 0.9513 0.7792 0.7850
identity 1 0.8597 0.8604 0.7985 0.7792 0.8056 0.8077 0.9291 0.9402 0.7658 0.7695
ResNet18 3 0.8968 0.8989 0.8976 0.8900 0.8071 0.8227 0.8974 0.9139 0.7920 0.7955
Table 6: Performance (AUC and F1) for different token mixers $T$ across datasets. The MetaFormer utilizes pooling at the first two stages and the given token mixer at the last two stages ($[P,P,T,T]$). These results provide the base for the calculation of the percental improvement shown in Fig. 4(b).
ImageWoof PathMNIST DermaMNIST PneumoniaMNIST OrganSMNIST
Token Mixer $T$ Kernel Size AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
pooling 3 0.9495 0.7907 0.9758 0.8767 0.8883 0.6868 0.9921 0.9656 0.9617 0.7689
5 0.9517 0.7965 0.9799 0.8773 0.8911 0.6855 0.9832 0.9446 0.9375 0.7576
7 0.9539 0.8059 0.9829 0.8986 0.8795 0.6991 0.9814 0.9537 0.9420 0.7706
conv 3 0.9466 0.7657 0.9697 0.8551 0.9035 0.7141 0.9755 0.9425 0.9594 0.7804
5 0.9191 0.7243 0.9785 0.8735 0.8456 0.6869 0.9841 0.9386 0.9611 0.7921
7 0.9036 0.6752 0.9862 0.9145 0.8526 0.6774 0.9864 0.9496 0.9671 0.7789
grouped conv 3 0.9357 0.7616 0.9708 0.8538 0.8865 0.7174 0.9894 0.9654 0.9468 0.7922
5 0.9208 0.7166 0.9826 0.9066 0.8793 0.6845 0.9895 0.9569 0.9474 0.7798
7 0.9089 0.6836 0.9846 0.8991 0.8607 0.7012 0.9857 0.9353 0.9506 0.7918
local attn 3 0.8796 0.5225 0.9771 0.8529 0.9168 0.5749 0.9716 0.9083 0.9702 0.7562
5 0.8691 0.4881 0.9690 0.8074 0.9179 0.5452 0.9734 0.9171 0.9721 0.7598
7 0.8620 0.4783 0.9733 0.8096 0.9230 0.5553 0.9665 0.9082 0.9701 0.7447
global attn - 0.8907 0.5486 0.9823 0.8470 0.9328 0.5944 0.9638 0.8582 0.9686 0.7646
identity 1 0.9659 0.7852 0.9560 0.7465 0.9474 0.7106 0.9892 0.9533 0.9637 0.7750
Table 7: Detailed results by our ranking schema (cf. Sec. 4.3) for each token mixer across datasets. Each entry shows a tuple: the first element is the absolute number of pairwise wins against other token mixers, and the second is the normalized rank in $[0.1,1]$. Global ranks are obtained as the geometric mean of normalized ranks across datasets.
Dataset ImageWoof Path Derma Pneumonia OrganS geometric mean
Token Mixer $T$ Kernel MedMNIST
$[T,T,T,T]$ from scratch
pooling 3 (8, 0.72) (2, 0.34) (2, 0.79) (5, 0.9) (11, 0.9) 0.691
5 (8, 0.72) (4, 0.55) (2, 0.79) (1, 0.41) (1, 0.2) 0.484
7 (9, 0.9) (8, 0.79) (1, 0.52) (1, 0.41) (2, 0.38) 0.563
conv 3 (8, 0.72) (9, 0.86) (0, 0.24) (1, 0.41) (8, 0.76) 0.541
5 (5, 0.48) (12, 0.97) (2, 0.79) (1, 0.41) (6, 0.58) 0.616
7 (4, 0.38) (12, 0.97) (0, 0.24) (7, 1.0) (7, 0.65) 0.563
grouped conv 3 (9, 0.9) (4, 0.55) (2, 0.79) (5, 0.9) (8, 0.76) 0.767
5 (7, 0.58) (6, 0.72) (1, 0.52) (1, 0.41) (2, 0.38) 0.508
7 (5, 0.48) (4, 0.55) (0, 0.24) (3, 0.76) (4, 0.52) 0.477
local attn 3 (2, 0.27) (0, 0.13) (2, 0.79) (1, 0.41) (2, 0.38) 0.34
5 (2, 0.27) (0, 0.13) (0, 0.24) (1, 0.41) (0, 0.1) 0.205
7 (1, 0.17) (2, 0.34) (0, 0.24) (1, 0.41) (1, 0.2) 0.259
global attn - (0, 0.1) (4, 0.55) (13, 1.0) (0, 0.1) (13, 1.0) 0.353
identity 1 (11, 1.0) (1, 0.24) (1, 0.52) (3, 0.76) (11, 0.9) 0.609
$[P,P,T,T]$ pretrained channel-MLPs
pooling $P$ 3 (2, 0.36) (1, 0.18) (0, 0.52) (0, 0.36) (6, 0.52) 0.366
5 (2, 0.36) (3, 0.36) (0, 0.52) (0, 0.36) (14, 0.95) 0.474
7 (6, 0.68) (8, 0.74) (0, 0.52) (0, 0.36) (1, 0.26) 0.477
conv 3 (3, 0.58) (1, 0.18) (0, 0.52) (0, 0.36) (6, 0.52) 0.401
5 (0, 0.15) (4, 0.42) (0, 0.52) (0, 0.36) (11, 0.89) 0.405
7 (0, 0.15) (7, 0.55) (0, 0.52) (0, 0.36) (7, 0.71) 0.409
grouped conv 3 (2, 0.36) (2, 0.29) (0, 0.52) (5, 0.95) (1, 0.26) 0.422
5 (0, 0.15) (2, 0.29) (0, 0.52) (4, 0.81) (6, 0.52) 0.396
7 (13, 0.81) (14, 0.89) (0, 0.52) (0, 0.36) (1, 0.26) 0.514
local attn 3 (3, 0.58) (8, 0.74) (0, 0.52) (0, 0.36) (1, 0.26) 0.462
5 (2, 0.36) (7, 0.55) (0, 0.52) (7, 1.0) (7, 0.71) 0.595
7 (3, 0.58) (7, 0.55) (0, 0.52) (0, 0.36) (7, 0.71) 0.533
local attn (warm start) 3 (10, 0.74) (8, 0.74) (0, 0.52) (0, 0.36) (8, 0.84) 0.613
5 (15, 0.95) (7, 0.55) (0, 0.52) (4, 0.81) (1, 0.26) 0.565
7 (15, 0.95) (14, 0.89) (0, 0.52) (0, 0.36) (15, 1.0) 0.695
global attn - (2, 0.36) (14, 0.89) (0, 0.52) (1, 0.68) (4, 0.42) 0.546
(warm start) - (15, 0.95) (17, 1.0) (0, 0.52) (4, 0.81) (0, 0.1) 0.526
identity 1 (13, 0.81) (0, 0.1) (17, 1.0) (4, 0.81) (7, 0.71) 0.543
Table 8: Detailed results by our ranking schema (cf. Sec. 4.3) for each token mixer across datasets. Each entry shows a tuple: the first element is the absolute number of pairwise wins against other token mixers, and the second is the normalized rank in $[0.1,1]$. Global ranks are obtained as the geometric mean of normalized ranks across datasets.
Dataset JSRT GRAZ Tiger@768²
TokenMixer $T$ Kernel geometric mean
$[T,T,T,T]$ from scratch
pooling 3 (4, 0.55) (2, 0.45) (6, 0.92) 0.608
5 (3, 0.45) (2, 0.45) (2, 0.55) 0.478
7 (1, 0.27) (2, 0.45) (2, 0.55) 0.406
9 (6, 0.92)
conv 3 (8, 0.69) (5, 0.72) (10, 1.0) 0.793
5 (10, 1.0) (6, 0.79) (1, 0.41) 0.687
7 (9, 0.86) (10, 1.0) (1, 0.41) 0.707
9 (3, 0.69)
grouped conv 3 (9, 0.86) (2, 0.45) (2, 0.55) 0.596
5 (8, 0.69) (9, 0.93) (5, 0.83) 0.811
7 (9, 0.86) (8, 0.86) (4, 0.77) 0.832
9 (3, 0.69)
local attn 3 (0, 0.1) (2, 0.45) (0, 0.21) 0.212
5 (4, 0.55) (2, 0.45) (0, 0.21) 0.374
7 (1, 0.27) (2, 0.45) (0, 0.21) 0.296
9 (0, 0.21)
global attn - (1, 0.27) (0, 0.13) DNS 0.192
identity 1 (1, 0.27) (0, 0.13) (0, 0.21) 0.198